Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $1.00 coming 3 years from today. The dividend should grow rapidly - at a rate of 36% per year - during Years 4 and 5; but after Year 5, growth should be a constant 6% per year. If the required return on Computech is 18%, what is the value of the stock today? Round your answer to the nearest cent. Do not round your intermediate calculations.
Answer #1

Year 3 dividend = 1

Year 4 dividend = 1 (1 + 36%) = 1.36

Year 5 dividend = 1.36 (1 + 36%) = 1.8496

Year 6 dividend = 1.8496 (1 + 6%) = 1.960576

Value at year 5 = D6 / required rate - growth rate

Value at year 5 = 1.960576 / 0.18 - 0.06

Value at year 5 = 1.960576 / 0.12

Value at year 5 = 16.338133

Value fo stock = 1 / (1 + 0.18)^3 + 1.36 / (1 + 0.18)^4 + 1.8496 / (1 + 0.18)^5 + 16.338133 / (1 + 0.18)^5

Value fo stock = 0.608631 + 0.701473 + 0.808477 + 7.141549

Value fo stock = $9.26